# Progressive-Resolution Policy Distillation: Leveraging Coarse-Resolution Simulation for Time-Efficient Fine-Resolution Policy Learning

### v1

@@ VI-D2 Results
VI-E Tradeoff between Time-Efficiency and Sample-Efficiency
VI-E1 Settings
As mentioned in Sec. II-C, various methods have enhanced RL’s sample efficiency, but less attention has been given to “time efficiency.” This section demonstrates that the proposed method outperforms previous sample-efficient approaches in terms of time efficiency. SAC [16] and its ensemble variant E-SAC [14] were evaluated within the PRPD and finest fixed-resolution with $\Delta=10$.
VI-E2 Results
As shown in Fig. 6(d), the proposed PRPD framework outperforms the fixed resolution $\Delta=10$ in terms of time efficiency across all RL methods. Ours (PPO) reached target success rate $\hat{\tau}$ at the finest resolution $\Delta=10$ in less than 100 minutes, significantly faster than the training in only finest resolution $\Delta=10$ (PPO), which took about 500 minutes. The extremely slow simulation at $\Delta=10$ made the $\Delta=10$ (PPO) less time-efficient than sample-efficient methods, as the sampling cost outweighed the policy optimization cost. Comparing ours (SAC) with $\Delta=10$ (SAC), there is little difference in sample numbers, but the learning time shows over a twofold difference, demonstrating that the progressive resolution shift significantly enhanced time efficiency.
VI-F Sim-to-Real Policy Transfer
This section validates the sim-to-real policy transfer across diverse real-world rock-excavation environments, including silica sand (low viscosity), kinetic sand (high viscosity), and plastic pellets (blocky shape and large grain size). Table II displays the evaluation of the transfer policies. The experiment shows that the success rate increases as the resolution improves in the real-world environment for fixed resolution policies, as was demonstrated with the simulation results. PRPD achieves performance comparable to fine-resolution fixed policy learning, indicating no adverse effects from using coarse-resolution simulations and successful sim-to-real transfer. We also evaluated the differences in policy performance during sim-to-real transfer depending on the presence of DR parameters. Table II shows that policy performance dropped to approximately 50 % without DR. This indicates a reality gap between the simulation and real-world environments, mitigated by applying DR. These results demonstrate that the variable-resolution rock excavation simulator and the proposed policy learning framework achieve approximately 90 % success rates for nine types of real rock environments.
@@ VII Discussions
Finer resolution assumption: In this paper, we assumed that finer resolution in excavation simulators leads to better transfer policy performance in real-world environments, supported by the experimental results in Sec. VI-F. However, this assumption may not always hold. The relationship between resolution and the reality gap is complex, as finer resolution may sometimes widen the gap. Moreover, when combined with domain randomization, simulators with relatively coarser resolution may enhance the learned policy’s performance and generalization.
Automatic determination of resolution interval: As an extension of the proposed PRPD, automatic determination of resolution intervals $\Delta_{\mathcal{R}}$ could be considered. In Sec. VI-D, we showed that finer $\Delta_{\mathcal{R}}$ allows for learning with fewer samples. Since these parameters depend on the task and policy learning algorithm, determining a single optimal value is difficult. However, developing a framework to automatically adjust simulation intervals based on their relationship would be beneficial.
Applicability of proposed framework: This paper focuses on rock excavation within earthwork tasks, developing a simulator and a learning algorithm. PRPD, which progressively changes simulation resolution, could extend to other excavation tasks or the installation of non-rock objects. It may also apply to tasks where simulation resolution impacts simulation time beyond earthwork. As shown in Fig. 5, this study achieved a 7-fold improvement in learning time efficiency, an approximately 8-hour time gap. This time gap may grow with increasing task complexity or state-action space. Using finer resolutions than $\Delta=10$ could further widen the learning time difference.
@@ VIII Conclusion
In this paper, we proposed time-efficient policy learning framework PRPD to address the time inefficiency of a fine-resolution rock-excavation simulator. We evaluated PRPD by developing a variable-resolution rock-excavation simulator using Isaac Gym. PRPD significantly reduced policy learning time in the simulator. Additionally, the learned policy was successfully transferred to the real-world environment, robustly removing previously unseen rocks.

### v2

@@ VII Discussions
Finer resolution assumption: In this paper, we assumed that finer resolution in excavation simulators leads to better transfer policy performance in real-world environments, supported by the experimental results in Section VI-F. However, this assumption may not always hold. The relationship between resolution and the reality gap could reverse or remain unchanged beyond a certain point. While finer resolutions may reveal physical differences, these might not significantly impact policy learning. The influence of policy generalization through DR on this relationship is also considered. These discussions are left for future work.
Automatic determination of resolution interval: As an extension of the proposed PRPD, automatic determination of resolution intervals $\Delta_{\mathcal{R}}$ could be considered. In Section VI-D, we showed that finer $\Delta_{\mathcal{R}}$ allows for learning with fewer samples. Since these parameters depend on the task and policy learning algorithm, determining a single optimal value is difficult. However, developing a framework to automatically adjust simulation intervals based on their relationship would be beneficial.
Applicability of proposed framework: This paper focuses on rock excavation within earthwork tasks in the development of a simulator and an effective learning algorithm. The proposed RL framework of progressively changing simulation resolution for faster policy learning could be extended to other excavation tasks or the installation of non-rock objects. Additionally, this approach could broadly apply to tasks where simulation resolution affects simulation time beyond just earthwork tasks. This paper demonstrated a 7-fold improvement in learning time efficiency (Ours: 90 minutes, fixed highest resolution: 600 minutes), resulting in an 8-hour time gap. This difference is likely to become more significant with increasing state-action space complexity or task difficulty. Furthermore, while the finest resolution in this study was set to $\Delta=10$, targeting even higher resolutions would likely widen the gap in learning time.
@@ VIII Conclusion
In this paper, we propose time-efficient RL framework PRPD to address the time inefficiency of a fine-resolution rock-excavation simulator. We evaluated PRPD by developing a variable-resolution rock-excavation simulator using Isaac Gym. PRPD significantly reduced policy learning time in the simulator. Additionally, the learned policy was successfully transferred to the real-world environment, robustly removing previously unseen rocks.